• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 23:25:58 on June12,2004

    Entry number 125494

    keyword=Owl Shift Summary - better late than never.

    Owl Shift Summary
    
    The shift began with a reset of Q1 which had tripped due to a dip in flow from the ESR.  At the same time, Dave Meekins was here to refill the target.  The hall went into controlled access in order to fix both of these problems.
    
    The reset of Q1 did not work, so Heidi was called in to fix it (she gave us phone support during our first and second attempts at resetting it).  She was able to bring Q1 back online, and she was out of the hall by 3 am.
    
    Dave worked on the target until ~3:30 am.  He decided to set the pressure of the target at 175 psi to perhaps alleviate some problems with the CHL flow.
    
    While we were having our troubles in the hall, the south linac of the accelerator was down.  Since Hall B was down, we asked if we could take some data in the injector.  We were able to take one RHWP scan (Run 2040) before the south linac was up, and beam was restored to Hall B.  While Hall A was still down, we took some HAPPEX runs to measure pedestals/electronics noise.
    
    Run 2041 - empty target
        2042 - empty target
    
    Once all the work on the magnets, and target were done, we put the 4He target back in the beamline and began taking data.  We decided to slowly increase the current so that we could watch the target and make sure it was ok along the way.
    
    Run 2043 - 4He, 175 psi, 2x2 raster, 10 uA at beginning and then up to 20 uA (left dipole momentum wrong - 2.87 not 3.009)
    Run 2044 - 4He, 175 psi, 2x2 raster, 20 uA at beginning and then up to 25 uA (maybe 30 uA)
    Run 2045 - started and stopped right away in order to turn off feedback to get ready for some RHWP scans
    Run 2046 - RHWP scan, IHWP out, PITA offset = 0. 4He, 2x2 raster, 29 uA
    Run 2047 - RHWP scan, IHWP out, PITA offset = -200, 4He, 2x2 raster, 29 uA
    Run 2048 - RHWP scan, IHWP in, PITA offset = 0, 4He, 2x2 raster, 29 uA
    Run 2049 - 4He, 2x2 raster, 29 uA, feedback on again
    
    Unfortunately, we found out at the end of our shift that the quadrupoles on the left arm had been set incorrectly during the night.  Please see logentries: 125439, 125467, and 125471.
    
    Also, during the shift, there were problems with the target gui.  All the screens went white.  It was thought that the target ioc should be rebooted, but the screens did not come back after the reboot.  We called JP while also reviewing the target manual.  It ended up that the target gui needed to be restarted.  Everything was ok with the target itself during this time.  See logentry 125421.